how many quarts of 20% saline solution must be mixed with some 50% saline solution to obtain a mixture of 30 quarts of mixture that is 40%?
Answer by nerdybill(5411)

Let x = quarts of 20% saline
then
30-x = quarts of 50% saline
.
Our equation:
.20x + .50(30-x) = .40(30)
.20x + 15 - .50x = 12
15 - .30x = 12
15 = .30x + 12
3 = .30x
10 quarts = x (amount of 20% saline solution)
.
50% saline solution:
30-x = 30-10 = 20 quarts (amount of 50% saline solution)